April 8, 2015 Calibration
GlueX Calibration Meeting
Wednesday, April 8, 2015
11:00 am, EDT
JLab: CEBAF Center, F326
Connection Using Bluejeans
- To join via Polycom room system go to the IP Address: 22.214.171.124 (bjn.vc) and enter the meeting ID: 630804895.
- To join via a Web Browser, go to the page  https://bluejeans.com/630804895.
- To join via phone, use one of the following numbers and the Conference ID: 630804895
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4
- More information on connecting to bluejeans is available.
- Subdetector Reports
- RF Signals (Paul) Logbook Entry
- Timing Updates
- TDC reconstruction (Paul) Mark, Kei Doc
- CCDB updates (Mike S.)
- Spring Calibration Plans
- Normalization Discussion
Attending: Sean (NU); Simon, Adesh, Nathan, Paul, Luke, Beni, Will M., Eugene, Mike S., Justin, Alex B. (JLab); Curtis, Naomi (CMU); Noemi (Regina)
- FCAL - Adesh has been working on trigger studies, and studying expectations for rates for the upcoming run. He truncated the photon spectrum at 6 GeV, and estimated the the integrated rates to be ~30-40% of the fall running (for same beam luminosity and running time). Higher beam currents and thicker radiators would be helpful for this running.
- BCAL - No major improvements to report. Tegan has been working on improving the simulation of dark hits in the BCAL simulation. George from Athens is working on determining the effective velocities.
- CDC - Beni has been working on trying to do an alignment with the cosmic data that's been taken over the past couple months. This corresponds to much more data than before, ~100M events. He is talking with Simon to optimize the processing of this data, since it's otherwise pretty slow. He is also working on tuning the gas flow (apparently the mass flow controller does not like Argon). Naomi is still working on the monitoring, but did find that when Beni turned off the alcohol bubbler the CDC gains when down ~20% and the maximum drift time increased by ~20 ns.
- FDC - Lubomir has finished aligning each of the packages by themselves, and it working on aligning the packages with each other. Simon is testing the implementation of these constants in sim-recon. Lubomir's alignment aligns the cathode strip planes with respect to the wire planes, and uses 48 parameters - 8 per cell - consisting offsets, rotations, and pitches.
- TOF - Kei is working on looking at the timing offsets, and Beni has been looking at the effective speeds of light. Sean is looking at the TOF/track matching. Sasha O. has been supposed to add his preliminary constants from cosmic data, but has not done this yet. Sean will continue to follow up on this.
- Start Counter - No news/
- TAGH - Nathan has been doing a lot of work with PS data. He has been trying to calibrate the coarse PS TDC times by cross-correlating paddles in the different arms. The offsets seem small (<1ns), and the resolution is ~250 ps. He has been using these PS calibrations to improve the hodoscope counter-by-counter calibrations. These calibrations limited by the PS energy range to a little below the microscope, and are statistically limited (esp. at lower photon energies). The resolutions obtained so far are ~250-400 ps.
- TAGM - No news.
Paul discussed some updates on the RF timing signals. First, he pointed to a logbook entry that discussed the installation of these signals. Paul has added these signals to the translation table and reconstruction code. He is working on a plugin to do monitoring and calibration of these signals.
Paul first pointed to a recently added GlueX note by Mark Ito and Kei that described the peculiarities of the F1TDC and CAEN TDC timing measurements and calibrations. Paul recently changed how the F1TDC times are handled in the reconstruction code. First, all times are referenced to the global system time provided by the TI. Second, conversion from TDC times to physical times (in ns) is being done in a consistent way, using the configuration information stored in the data stream since late in the Fall, when available, instead of using CCDB constants. However, Kei just found out that this information has been incorrect, and is working with David L. to fix this. So now, all taken already taken will use CCDB data, and the new scheme will be used going forward once this issue is fixed. A hardcoded cut on the run number is used to differentiate the epochs.
All this changed the timing calibrations Mike Staib had been working on.
Mike has run his timing calibration code over the fall data 3 times:
- Using the best, high luminosity runs. Unfortunately, this did not cover all the changes in timing offsets that were made (e.g. the configurations for modes 7 and 8 were often subtly different).
- Using looser requirements. This worked well.
- After Paul's TDC timing changes that were described above.
- Once more is needed after Kei's discovery. Luckily, this running is all automated now.
His code does several calibration steps:
- Aligns overall ADC/TDC times.
- Aligns TDCs for individual detector elements with the ADC times.
- Aligns the tagger times with respect to the PS times.
The results from these could also be used to investigate some more detailed per-channel timing calibrations. Sean will look at the results for the TOF.
Spring Calibration Plans
We reviewed the updated plans, and no one had any comments. In regards to normalization, Simon and Kei still need to meet with Sascha, and Sean and Dave Mack have been discussing plans for measuring QED processes. More study of these topics is needed.
Next meeting will include a more detailed discussion of plans for measuring the performance of the detector.